# The Quest for the Right Mediator: Surveying Mechanistic Interpretability Through the Lens of Causal Mediation Analysis

### v2

@@ Abstract
Interpretability provides a toolset for understanding how and why language models behave in certain ways. However, there is little unity in the field: most studies employ ad-hoc evaluations and do not share theoretical foundations, making it difficult to measure progress and compare the pros and cons of different techniques. Furthermore, while mechanistic understanding is frequently discussed, the basic causal units underlying these mechanisms are often not explicitly defined. In this article, we propose a perspective on interpretability research grounded in causal mediation analysis. Specifically, we describe the history and current state of interpretability taxonomized according to the types of causal units (mediators) employed, as well as methods used to search over mediators. We discuss the pros and cons of each mediator, providing insights as to when particular kinds of mediators and search methods are most appropriate. We argue that this framing yields a more cohesive narrative of the field and helps researchers select appropriate methods based on their research objective. Our analysis yields actionable recommendations for future work, including the discovery of new mediators and the development of standardized evaluations tailored to these goals.
@@ 1 Introduction
To understand how language models (LMs) generalize, we must understand the causes of their behavior. These causes include inputs, but also the intermediate computations of the network; this survey is concerned with understanding these intermediate computations. How can we understand what an LM’s computations represent, such that we can obtain a deeper algorithmic understanding of how and why models behave the way they do? For example, if a model decides to refuse a user’s request, was the refusal mediated by an underlying concept of toxicity, by the presence of superficial correlates of toxicity (such as the mention of particular demographic groups), or some other unexpected variable? The former would be significantly more likely to robustly and safely generalize. These questions motivate the field of mechanistic interpretability (MI), which aims to understand how LMs arrive at particular behaviors by understanding the functional roles of their components.
@@
Figure 1: Outline of survey. We first define necessary causal terminology (§2) and contextualize our perspective with others’ (§3). We then give an overview of the history of mechanistic interpretability centered on units of causal analysis (§4). We then survey and categorize commonly used units of analysis and describe their strengths and weaknesses (§5), as well as methods for searching over them (§6). Finally, we discuss (§7) what we consider to be among the most important questions in mechanistic interpretability: What are the right causal abstractions for understanding and discussing the inner workings of LMs7.1)? What kinds of mediators and research will be needed to advance the field7.2)?
@@
We view mechanistic interpretability [^1] The meaning of “mechanistic interpretability” is debated; see Saphra and Wiegreffe (2024). We define the term as any study that aims to understand, explain, or modify a neural network’s behavior by studying the model’s internal components, such as its representations or weights. as equivalent to extracting causal graphs explaining how intermediate LM computations mediate model outputs. This framing based in causality enables a new perspective of the field: prior surveys have organized the field according to methodological differences, whereas we taxonomize work in the field according to the kinds of causal mediators—or types of nodes in the causal graphs—that a study employs (e.g., neurons, non-basis-aligned directions, attention heads, and the like). We start by describing causal mediation analysis and its role in MI (§2); we also define the goals of MI, and goal-specific criteria by which the success of a mediator can be measured (§2.1). We then contrast this survey with other MI surveys (§3), noting in particular the lack of surveys that center the mediator type. Following this, we present a history of mechanistic interpretability for neural networks more broadly (§4), from backpropagation to the beginning of the current wave of mechanistic interpretability research.
@@
Finally, after surveying the field, we discuss practical considerations and implications for future work (§7). We point out mediators which have been underexplored, but have significant potential to yield new insights; propose future mediators that are likely satisfy the criteria laid out in §2.1; and suggest ways to measure progress in mechanistic interpretability moving forward. Figure 1 summarizes the content of this survey. [^2] Note that the methods discussed in this survey generalize beyond language models: most can, in theory, apply to any neural network. Our focus on language models is motivated by (i) the centrality of mechanistic interpretability in natural language processing research in recent years, (ii) our primary expertise lying in this field, and (iii) a desire to keep the survey size tractable.
@@ 2 Preliminaries
where “do” is an operation where the experimenter intervenes, setting a value in the experiment to one that it would not naturally have taken without the experimenter’s involvement. See Figure 2 for an illustration of the indirect effect given a language model component ZZ. [^8] The choice of counterfactual $v_{i}^{\prime}$ significantly affects which components will be identified as influential. Common approaches for selecting $v_{i}^{\prime}$ include using values from alternative inputs $x^{\prime}$, applying a constant (e.g., 0), or adding noise $\epsilon$, where $\epsilon$ is typically scaled according to the norm or variance of $v_{i}$. These are conceptually distinct operations that will uncover different kinds of components; we provide a more detailed discussion in Appendix A. We leave this to the appendix, as this survey is concerned with how results are affected by the choice of mediator ZZ—a topic distinct from the choice of intervention value $v_{i}^{\prime}$.
@@ 4 Lessons from the History of Interpretability
Hence, correlational probing methods rose to meet this demand. In this approach, classifiers are trained on intermediate activations to extract some target phenomenon. Probing classifiers have been used to investigate the latent morphosyntactic structures encoded in static word embeddings (Köhn, 2015; Gupta et al., 2015) or intermediate hidden representations in pre-trained language models—for example, in neural machine translation systems (Shi, Padhi, and Knight, 2016; Belinkov et al., 2017; Conneau et al., 2018) and pre-trained language models (Hewitt and Manning, 2019; Hewitt et al., 2021; Lakretz et al., 2019, 2021). However, probing classifiers lack consistent baselines, and the claims made in these studies were not often causally verified (Belinkov, 2022). For instance, although an intervention may target a causal property of the task $V_{i}\rightarrow Y$, an alternative spurious property $V_{j}$ may be picked up by the probe, which impedes causal claims about $V_{i}\rightarrow Y$ (Ravichander, Belinkov, and Hovy, 2021). This encouraged researchers to search for more causally efficacious methods.
@@ 5 Mediator Types
In this section, we discuss different types of causal mediators in neural networks, and the pros and cons of each. Figure 3 visualizes a computation graph of a Transformer block in a typical language model, and units in the graph that are often used as mediators in mechanistic interpretability studies. In mechanistic interpretability, we often do not want to treat the full computation graph as the final causal graph, as it is large and difficult to directly interpret. Thus, we typically want to build higher-level causal abstractions that capture only the most important mediators, and/or where each causal node is human-interpretable.
@@ 5.2 Basis-aligned subspaces
Similar to neurons, attention heads are fundamental components of Transformer-based neural networks. They mediate the flow of information between token positions (Vaswani et al., 2017); thus, using attention heads as units of causal analysis can help us understand how models synthesize contextual information (Ma et al., 2021; Neo, Cohen, and Barez, 2024) to predict subsequent tokens (Wang et al., 2023; Hanna, Liu, and Variengien, 2023b; Prakash et al., 2024; Garc´ıa-Carrasco, Maté, and Trujillo, 2024; Brinkmann et al., 2024).
@@
Attention heads are attractive mediators because they are easily enumerable: there are generally far fewer attention heads $O(\ell\cdot|A^{\ell}|)$than neurons $O(\ell\cdot d)$ in a model, as $|A^{\ell}|\ll d$ in typical language models. Attention heads also track multi-token relationships. However, in contrast to the activation $h_{i}^{\ell}$ of a neuron, the output of an attention head $\mathbf{a}_{i}^{\ell}$ is multidimensional. Thus, it is difficult to directly interpret the full set of functional roles a single head might have: attention heads are almost always polysemantic, so one cannot typically determine the function(s) of an attention head solely by observing its outputs (Janiak, cmathw, and Heimersheim, 2023). [^21] However, there is initial evidence that some dimensions of an attention head’s output can be meaningfully explained (Merullo, Eickhoff, and Pavlick, 2024a, b; Hu et al., 2025). Thus, by decomposing the vector output of a head into smaller subspaces or even individual neurons, it may be easier to explain the set of functional roles of a given head. It has additionally been observed that intervening on an attention head can cause other attention heads to compensate, which further complicates causal analyses (Jermyn, Olah, and Henighan, 2023; Wang et al., 2023; McGrath et al., 2023). [^22] This phenomenon where downstream components only have causal relevance after an upstream component has been removed is sometimes called preemption in the causality literature (Mueller, 2024), or the “Hydra effect” in mechanistic interpretability (McGrath et al., 2023). Preemption is not limited to attention heads; future work should analyze how common preemption is between other types of components, such as MLP submodules. To summarize, attention heads are easier to exhaustively search over than (sets of) neurons, but have the same issues of low selectivity and human-interpretability.
@@ 6.2.2 Unsupervised mediator searches
In summary, unsupervised mediator searches enable us to locate human-interpretable concepts in language models without any labeled data. They allow us to do so relatively quickly, and these concepts can then be causally implicated in model behavior post hoc. However, the recovered concepts are not guaranteed to be faithful to the model’s true concept space, and we are never guaranteed to recover a complete nor non-redundant concept set. Another pressing issue in unsupervised search is evaluation: while it is possible to compare the relative quality of two different unsupervised interpretability searches, it is difficult to devise absolute metrics that meaningfully capture closeness to some ideal solution; indeed, an ideal or identifiable solution may not exist (Méloux et al., 2025). For SAEs, some have tried to devise thorough evaluations Karvonen et al. (2025), but robust evaluations that enable comparisons between supervised and unsupervised methods have only just started to become common (e.g., Wu et al., 2025; Mueller et al., 2025).
@@ 7 Discussion
What is the right unit of analysis for describing the inner workings of language models? Thus far, we have categorized past work by mediator type and search method. Now, we turn our attention to practical questions: What kinds of studies can more easily be done with certain mediator types? Where is there still room to deploy less common mediators in useful ways? Our main goal is to encourage researchers to conceptualize the field in this way, in the hope that we can encourage more work on discovering better causal abstractions and better terms for discussing the inner workings of language models—and, more ambitiously, more rigorous theoretical foundations for interpretability.
@@ 7.1 What is the right mediator?
If we already have a hypothesis as to how a model performs some behavior and wish to measure the accuracy of our hypothesis, then multidimensional non-basis-aligned subspaces5.3) may be the right mediators, and a reasonable corresponding search method would be counterfactual-based optimization6.2.1). One can automatically search for the subspaces which correspond to a particular node in one’s hypothesized causal graph using alignment search methods (Geiger et al., 2024; Wu et al., 2023; Sun et al., 2025; Geiger et al., 2021). Alignment search entails learning a linear transformation to isolate some target concept; this allows us to locate distributed representations that act as independent causal variables in non-basis-aligned spaces. This is relatively scalable, and enables us to qualitatively understand intermediate model computations. The primary downside is that we must anticipate the mechanisms that models employ to perform a task, as demonstrated in Appendix C; if we cannot anticipate them, then curating data and refining one’s causal hypotheses may require significant human effort. If these are significant concerns, then probing may be a better (though correlative and less precise) option. These methods are subject to the same confounds as any other optimization-based technique: the module we are optimizing could directly learn the phenomenon, rather than extracting it from the language model (Hewitt and Liang, 2019; Sutter et al., 2025).
@@ 7.2.1 Finding better causal mediators
As discussed in §5.4, there is recent work demonstrating the existence of human-interpretable multidimensional features. For example, days of the week are encoded circularly as a set of 7 directions in a two-dimensional subspace (Engels et al., 2025), and current methods cannot easily capture these multidimensional features. The ability to discover these features could greatly improve our understanding of the feature space of language models, and thus our ability to systematically explain more of their behavior (i.e., increase generality) in a faithful and human-interpretable way. Group sparse autoencoders (Theodosis and Ba, 2023) or clusters of autoencoder features could be a way to capture multidimensional non-basis-aligned features in an unsupervised manner, but despite promising initial evidence, empirical work has not yet demonstrated to what degree this will be effective for interpreting or controlling language models. Additionally, many current mechanistic interpretability methods require us to define binary distinctions between correct and incorrect answers, whereas causal mediation analysis does not have any theoretical linearity, dimensionality, or Boolean restrictions; thus, interpretability methods will need to be extended to handle new kinds of variables.
@@
There may also exist higher-order non-linear concepts in latent space that we have not yet located due to the linear focus of contemporary methods. For example, a subgraph or subcircuit can encode a coherent variable representation or functional role, as in Lepori, Serre, and Pavlick (2024); Li, Davies, and Nadeau (2024). How can we discover these subgraphs? Path patching (Goldowsky-Dill et al., 2023; Wang et al., 2023) provides a manual approach to implicating subgraphs as causal mediators, but we do not yet have automatic methods that can scalably search over subgraphs of a computation graph. Even if we were able to locate them, how might non-linear and/or coarse-grained mediators like these be useful in practice? As an example, we might expect fundamental phenomena like syntax processing to be spread across many layers of a model. Syntax-sensitive components should be implicated in downstream tasks like question answering (QA) if we expect that language models are robustly parsing the meaning of the inputs. Thus, one could use causal mediation analysis to locate all components in the model with some high indirect effect on syntactic processing (e.g., using a subject–verb agreement task); this can be conceptualized as the syntax processing region of the model. Then, one could implicate the region(s) in the model’s QA performance by intervening on each component in the region and observing whether performance changes. If the syntax processing region is not strongly implicated in QA performance, then we have a strong hint that the model may not be parsing the meaning of the inputs, but instead relying on a mixture of surface-level spurious heuristics—for example, memorized bigram associations, or giving answers with high prior probabilities. These examples illustrate how these more coarse-grained mediators could help us verify new kinds of mechanistic hypotheses.

### v3

@@ Abstract
Interpretability provides a toolset for understanding how and why neural networks behave in certain ways. However, there is little unity in the field: most studies employ ad-hoc evaluations and do not share theoretical foundations, making it difficult to measure progress and compare the pros and cons of different techniques. Furthermore, while mechanistic understanding is frequently discussed, the basic causal units underlying these mechanisms are often not explicitly defined. In this article, we propose a perspective on interpretability research grounded in causal mediation analysis. Specifically, we describe the history and current state of interpretability taxonomized according to the types of causal units (mediators) employed, as well as methods used to search over mediators. We discuss the pros and cons of each mediator, providing insights as to when particular kinds of mediators and search methods are most appropriate. We argue that this framing yields a more cohesive narrative of the field and helps researchers select appropriate methods based on their research objective. Our analysis yields actionable recommendations for future work, including the discovery of new mediators and the development of standardized evaluations tailored to these goals.
@@ 1 Introduction
To understand how a neural network (NN) will generalize, we must understand the causes of its behavior. These causes include inputs, but also the intermediate computations of the network; this survey is concerned with understanding these intermediate computations. How can we understand what a NN’s computations represent, such that we can obtain a deeper algorithmic understanding of how and why models behave the way they do? For example, if a language model (LM) decides to refuse a user’s request, was the refusal mediated by an underlying concept of toxicity, by the presence of superficial correlates of toxicity (such as the mention of particular demographic groups), or some other unexpected variable? The former would be significantly more likely to robustly and safely generalize. These questions motivate the field of mechanistic interpretability (MI), which aims to understand how NNs arrive at particular behaviors by understanding the functional roles of their components.
@@
Figure 1: Outline of survey. We first define necessary causal terminology (§2) and contextualize our perspective with others’ (§3). We then give an overview of the history of mechanistic interpretability centered on units of causal analysis (§4). We then survey and categorize commonly used units of analysis and describe their strengths and weaknesses (§5), as well as methods for searching over them (§6). Finally, we discuss (§7) what we consider to be among the most important questions in mechanistic interpretability: What are the right causal abstractions for understanding and discussing the inner workings of NNs7.1)? What kinds of mediators and research will be needed to advance the field7.2)?
@@
We view mechanistic interpretability [^1] The meaning of “mechanistic interpretability” is debated; see Saphra and Wiegreffe (2024). We define the term as any study that aims to understand, explain, or modify a neural network’s behavior by studying the model’s internal components, such as its representations or weights. as equivalent to extracting causal graphs explaining how intermediate NN computations mediate model outputs. This framing based in causality enables a new perspective of the field: prior surveys have organized the field according to methodological differences, whereas we taxonomize work in the field according to the kinds of causal mediators—or types of nodes in the causal graphs—that a study employs (e.g., neurons, non-basis-aligned directions, attention heads, and the like). We start by describing causal mediation analysis and its role in MI (§2); we also define the goals of MI, and goal-specific criteria by which the success of a mediator can be measured (§2.1). We then contrast this survey with other MI surveys (§3), noting in particular the lack of surveys that center the mediator type. Following this, we present a history of mechanistic interpretability for neural networks more broadly (§4), from backpropagation to the beginning of the current wave of mechanistic interpretability research.
@@
Finally, after surveying the field, we discuss practical considerations and implications for future work (§7). We point out mediators which have been underexplored, but have significant potential to yield new insights; propose future mediators that are likely satisfy the criteria laid out in §2.1; and suggest ways to measure progress in mechanistic interpretability moving forward. Figure 1 summarizes the content of this survey. [^2] Note that the methods discussed in this survey focus primarily on language models, so we will often use “language model” in place of “neural network” throughout this survey. However, most of the methods we discuss generalize beyond language models: most can, in theory, apply to any neural network.
@@ 2 Preliminaries
where “do” is an operation where the experimenter intervenes, setting a value in the experiment to one that it would not naturally have taken without the experimenter’s involvement. See Figure 2 for an illustration of the indirect effect given a neural network component ZZ. [^8] The choice of counterfactual $v_{i}^{\prime}$ significantly affects which components will be identified as influential. Common approaches for selecting $v_{i}^{\prime}$ include using values from alternative inputs $x^{\prime}$, applying a constant (e.g., 0), or adding noise $\epsilon$, where $\epsilon$ is typically scaled according to the norm or variance of $v_{i}$. These are conceptually distinct operations that will uncover different kinds of components; we provide a more detailed discussion in Appendix A. We leave this to the appendix, as this survey is concerned with how results are affected by the choice of mediator ZZ—a topic distinct from the choice of intervention value $v_{i}^{\prime}$.
@@ 4 Lessons from the History of Interpretability
Hence, correlational probing methods rose to meet this demand. In this approach, classifiers are trained on intermediate activations to extract some target phenomenon. Probing classifiers have been used to investigate the latent morphosyntactic structures encoded in static word embeddings (Köhn, 2015; Gupta et al., 2015) or intermediate hidden representations in natural language systems—for example, in neural machine translation systems (Shi, Padhi, and Knight, 2016; Belinkov et al., 2017; Conneau et al., 2018) and pre-trained language models (Hewitt and Manning, 2019; Hewitt et al., 2021; Lakretz et al., 2019, 2021). However, probing classifiers lack consistent baselines, and the claims made in these studies were not often causally verified (Belinkov, 2022). For instance, although an intervention may target a causal property of the task $V_{i}\rightarrow Y$, an alternative spurious property $V_{j}$ may be picked up by the probe, which impedes causal claims about $V_{i}\rightarrow Y$ (Ravichander, Belinkov, and Hovy, 2021). This encouraged researchers to search for more causally efficacious methods.
@@ 5 Mediator Types
In this section, we discuss different types of causal mediators in neural networks, and the pros and cons of each. Figure 3 visualizes a computation graph of a Transformer-based language model, and units in the graph that are often used as mediators in mechanistic interpretability studies. In mechanistic interpretability, we often do not want to treat the full computation graph as the final causal graph, as it is large and difficult to directly interpret. Thus, we typically want to build higher-level causal abstractions that capture only the most important mediators, and/or where each causal node is human-interpretable.
@@ 5.2 Basis-aligned subspaces
Similar to neurons, attention heads are fundamental components of Transformer-based neural networks. They mediate the flow of information between token positions (Vaswani et al., 2017); thus, using attention heads as units of causal analysis can help us understand how models synthesize contextual information (Ma et al., 2021; Neo, Cohen, and Barez, 2024) to predict subsequent tokens (Wang et al., 2023; Hanna, Liu, and Variengien, 2023b; Prakash et al., 2024; García-Carrasco, Maté, and Trujillo, 2024; Brinkmann et al., 2024).
@@
Attention heads are attractive mediators because they are easily enumerable: there are generally far fewer attention heads $O(\ell\cdot|A^{\ell}|)$than neurons $O(\ell\cdot d)$ in a model, as $|A^{\ell}|\ll d$ in typical Transformer-based models. Attention heads also track multi-token relationships. However, in contrast to the activation $h_{i}^{\ell}$ of a neuron, the output of an attention head $\mathbf{a}_{i}^{\ell}$ is multidimensional. Thus, it is difficult to directly interpret the full set of functional roles a single head might have: attention heads are almost always polysemantic, so one cannot typically determine the function(s) of an attention head solely by observing its outputs (Janiak, cmathw, and Heimersheim, 2023). [^21] However, there is initial evidence that some dimensions of an attention head’s output can be meaningfully explained (Merullo, Eickhoff, and Pavlick, 2024a, b; Hu et al., 2025). Thus, by decomposing the vector output of a head into smaller subspaces or even individual neurons, it may be easier to explain the set of functional roles of a given head. It has additionally been observed that intervening on an attention head can cause other attention heads to compensate, which further complicates causal analyses (Jermyn, Olah, and Henighan, 2023; Wang et al., 2023; McGrath et al., 2023). [^22] This phenomenon where downstream components only have causal relevance after an upstream component has been removed is sometimes called preemption in the causality literature (Mueller, 2024), or the “Hydra effect” in mechanistic interpretability (McGrath et al., 2023). Preemption is not limited to attention heads; future work should analyze how common preemption is between other types of components, such as MLP submodules. To summarize, attention heads are easier to exhaustively search over than (sets of) neurons, but have the same issues of low selectivity and human-interpretability.
@@ 6.2.2 Unsupervised mediator searches
In summary, unsupervised mediator searches enable us to locate human-interpretable concepts from neural networks internals without any labeled data. They allow us to do so relatively quickly, and these concepts can then be causally implicated in model behavior post hoc. However, the recovered concepts are not guaranteed to be faithful to the model’s true concept space, and we are never guaranteed to recover a complete nor non-redundant concept set. Another pressing issue in unsupervised search is evaluation: while it is possible to compare the relative quality of two different unsupervised interpretability searches, it is difficult to devise absolute metrics that meaningfully capture closeness to some ideal solution; indeed, an ideal or identifiable solution may not exist (Méloux et al., 2025). For SAEs, some have tried to devise thorough evaluations Karvonen et al. (2025), but robust evaluations that enable comparisons between supervised and unsupervised methods have only just started to become common (e.g., Wu et al., 2025; Mueller et al., 2025).
@@ 7 Discussion
What is the right unit of analysis for describing the inner workings of neural networks? Thus far, we have categorized past work by mediator type and search method. Now, we turn our attention to practical questions: What kinds of studies can more easily be done with certain mediator types? Where is there still room to deploy less common mediators in useful ways? Our main goal is to encourage researchers to conceptualize the field in this way, in the hope that we can encourage more work on discovering better causal abstractions and better terms for discussing the inner workings of neural machine learning systems—and, more ambitiously, more rigorous theoretical foundations for interpretability.
@@ 7.1 What is the right mediator?
If we already have a hypothesis as to how a model performs some behavior and wish to measure the accuracy of our hypothesis, then multidimensional non-basis-aligned subspaces5.3) may be the right mediators, and a reasonable corresponding search method would be counterfactual-based optimization6.2.1). One can automatically search for the subspaces which correspond to a particular node in one’s hypothesized causal graph using alignment search methods (Geiger et al., 2024; Wu et al., 2023; Sun et al., 2025; Geiger et al., 2021). Alignment search entails learning a linear transformation to isolate some target concept; this allows us to locate distributed representations that act as independent causal variables in non-basis-aligned spaces. This is relatively scalable, and enables us to qualitatively understand intermediate model computations. The primary downside is that we must anticipate the mechanisms that models employ to perform a task, as demonstrated in Appendix C; if we cannot anticipate them, then curating data and refining one’s causal hypotheses may require significant human effort. If these are significant concerns, then probing may be a better (though correlative and less precise) option. These methods are subject to the same confounds as any other optimization-based technique: the module we are optimizing could directly learn the phenomenon, rather than extracting it from the model (Hewitt and Liang, 2019; Sutter et al., 2025).
@@ 7.2.1 Finding better causal mediators
As discussed in §5.4, there is recent work demonstrating the existence of human-interpretable multidimensional features. For example, days of the week are encoded circularly as a set of 7 directions in a two-dimensional subspace (Engels et al., 2025), and current methods cannot easily capture these multidimensional features. The ability to discover these features could greatly improve our understanding of the feature space of NNs, and thus our ability to systematically explain more of their behavior (i.e., increase generality) in a faithful and human-interpretable way. Group sparse autoencoders (Theodosis and Ba, 2023) or clusters of autoencoder features could be a way to capture multidimensional non-basis-aligned features in an unsupervised manner, but despite promising initial evidence, empirical work has not yet demonstrated to what degree this will be effective for interpreting or controlling NNs. Additionally, many current mechanistic interpretability methods require us to define binary distinctions between correct and incorrect answers, whereas causal mediation analysis does not have any theoretical linearity, dimensionality, or Boolean restrictions; thus, interpretability methods will need to be extended to handle new kinds of variables.
@@
There may also exist higher-order non-linear concepts in latent space that we have not yet located due to the linear focus of contemporary methods. For example, a subgraph or subcircuit can encode a coherent variable representation or functional role, as in Lepori, Serre, and Pavlick (2024); Li, Davies, and Nadeau (2024). How can we discover these subgraphs? Path patching (Goldowsky-Dill et al., 2023; Wang et al., 2023) provides a manual approach to implicating subgraphs as causal mediators, but we do not yet have automatic methods that can scalably search over subgraphs of a computation graph. Even if we were able to locate them, how might non-linear and/or coarse-grained mediators like these be useful in practice? As an example, we might expect fundamental phenomena like syntax processing to be spread across many layers of a language model. Syntax-sensitive components should be implicated in downstream tasks like question answering (QA) if we expect that language models are robustly parsing the meaning of the inputs. Thus, one could use causal mediation analysis to locate all components in the model with some high indirect effect on syntactic processing (e.g., using a subject–verb agreement task); this can be conceptualized as the syntax processing region of the model. Then, one could implicate the region(s) in the model’s QA performance by intervening on each component in the region and observing whether performance changes. If the syntax processing region is not strongly implicated in QA performance, then we have a strong hint that the model may not be parsing the meaning of the inputs, but instead relying on a mixture of surface-level spurious heuristics—for example, memorized bigram associations, or giving answers with high prior probabilities. These examples illustrate how these more coarse-grained mediators could help us verify new kinds of mechanistic hypotheses.